Which of the following combinations represents the vertex and two x-intercepts of the function given below? y=x2-2x-24 O A. Vertex: (-1,-21); Intercepts: x = 6,4 O B. Vertex: (7,5); Intercepts. x = 6,8 o c. vertex. (1,-25): Intercepts: x = 6,4 O D. Vertex: (0,0); Intercepts: x =-4, 6
Answer #1

Solution :

\mathrm{C.\quad\:Vertex:}\:(1,-25);\:\quad\mathrm{intercepts:\:}\:x=6,\:-4

\mathrm{Description\::}

y=x^2-2x-24

a=1,\:b=-2

x=\frac{-\left(-2\right)}{2*1}

\mathrm{Simplify}

x=1

\mathrm{Plug\:in}\:x=1\:\mathrm{to\:find\:the}\:y\mathrm{\:value}

y=1^2-2*1-24

\mathrm{Simplify}

y=-25

\mathrm{Therefore\:the\:vertex\:is}\quad\left(1,\:-25\right)

\mathrm{and}

x-\mathrm{axis\:interception\:points\:of\:}x^2-2x-24:

x^2-2x-24=0

\mathrm{Solve\:with\:the\:quadratic\:formula}

\mathrm{For\:a\:quadratic\:equation\:of\:the\:form\:}ax^2+bx+c=0\mathrm{\:the\:solutions\:are\:}

x_{1,\:2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}

\mathrm{For\:}\quad a=1,\:b=-2,\:c=-24:\quad x_{1,\:2}=\frac{-\left(-2\right)\pm \sqrt{\left(-2\right)^2-4*1\left(-24\right)}}{2*1}

x_{1}=\frac{-\left(-2\right)+\sqrt{\left(-2\right)^2-4*1\left(-24\right)}}{2*1}

=\frac{-\left(-2\right)+\sqrt{\left(-2\right)^2-4*1\left(-24\right)}}{2}

=\frac{-\left(-2\right)+\sqrt{4+96}}{2}

=\frac{2+\sqrt{100}}{2}

=\frac{2+10}{2}

=\frac{12}{2}

=6

\Rightarrow x_{1}=6

\mathrm{and}

x_{2}=\frac{-\left(-2\right)-\sqrt{\left(-2\right)^2-4*1\left(-24\right)}}{2*1}

=\frac{-\left(-2\right)-\sqrt{\left(-2\right)^2-4*1\left(-24\right)}}{2}

=\frac{-\left(-2\right)-\sqrt{4+96}}{2}

=\frac{2-\sqrt{100}}{2}

=\frac{2-10}{2}

=\frac{-8}{2}

=-4

\Rightarrow x_{2}=-4

\mathrm{The\:final\:solutions\:to\:the\:quadratic\:equation\:are:}

x=6,\:x=-4

\mathrm{Therefore,\quad\:Intercepts}:\:x=6,\:-4